TIP: Click on subject to list as thread! ANSI
echo: astronomy
to: sci.space.news
from: baalke
date: 2011-05-06 22:00:40
subject: Cassini Update - May 6, 2011

From Newsgroup: sci.space.news


Cassini Significant Events 
for 04/27/11 - 05/03/11

The most recent spacecraft telemetry was acquired on May 3 from the 
Deep Space Network tracking complex at Canberra, Australia.  The 
Cassini spacecraft is in an excellent state of health and all 
subsystems are operating normally. Information on the present 
position and speed of the Cassini spacecraft may be found on the 
"Present Position" page at: 

http://saturn.jpl.nasa.gov/mission/presentposition/ .

Wednesday, April 27 (DOY 117)

Orbit Trim Maneuver (OTM) #282, an apoapsis maneuver scheduled for 
Apr. 29, was cancelled today.  A two second  bias in the time of 
closest approach at Titan 76 would have been required in order to 
make OTM-282 large enough to be executable.  By waiting until 
OTM-283, scheduled for May 5, to make the target correction, the 
required time bias is smaller and the orbit determination 
uncertainties decrease, leading to a more accurate flyby. The delta-V 
cost of canceling OTM 282 was very small.

Port 3 products were due today as part of the S69 Sequence 
Implementation Process (SIP).  The products will be merged and sent 
out to the flight team for review.

The Instrument Operations System (IOS)/Multi-Mission Image Processing 
Laboratory (MIPL) Cassini version 39.1 software delivery was approved 
to go operational at the Delivery Coordination Meeting (DCM) today. 
This delivery included multi-mission updates to the password server, 
enhancements and new capabilities for the Cassini Solstice Mission 
(CSM), bug fixes, and documentation updates.  Transition to 
operations is scheduled for May 11.

A DCM for Acelog V1.0, the replacement software tool for the 
operations log, was held this week.  Installation of Acelog into 
operations will occur within the next couple of weeks.

The Downlink Ground System (DGS) team resolved all major Java 
Distributed Object Manager (DOM) issues.  Java DOM Testing with the 
legacy Andrew File System (AFS) volumes will occur within the next 
three weeks prior to transitioning the DOM fully to the Network File 
System (NFS) next month.

Thursday, April 28 (DOY 118)

Dozens of John Muir High School students in the Business, Arts and 
Media, and Engineering and Environmental Science Academies were 
treated to an informative, entertaining, and educational field trip 
to JPL today where they individually shadowed a JPL scientist or 
engineer, toured the campus, and learned about preparing for college. 
Several Cassini team members participated in this event, taking their 
students to meetings and exposing them to work performed at JPL.  For 
more information on this event, link to: 

http://altadena.patch.com/articles/muir-high-students-check-out-jet-propulsion-laboratory
.

Friday, April 29 (DOY 120)

The Cassini Magnetospheric and Plasma Science (MAPS) workshop was 
held April 27-29 in Annapolis, MD.   Approximately 90 scientists 
attended, reporting results on topics including Saturn's rotational 
modulation, Saturn's aurorae, magnetospheric interactions with icy 
satellites including the Enceladus plume and Titan, and 
magnetospheric dynamics.

Port 1 products were due today as part of the S70 SIP.  The products 
will be merged and sent out to the flight team for review on May 2.

Saturday, April 30 (DOY 121)

Today at 2011-120T06:14:51, the Cassini power bus high rail shifted 
from a voltage of 20.7 V to 28.8 V and the bus low rail shifted from 
-8.9 V to -0.5 V.  A voltage reading of zero on the low rail would 
indicate that it was directly shorted to the chassis.  The 
Radioisotope Thermoelectric Generator (RTG) case voltages also 
shifted at that time - RTG1: 8.0 to 7.5 V, RTG2: 5.9 to 2.6 V, RTG3: 
7.9 to 1.3 V.  This is very similar to an incident that occurred on 
Aug. 18, 2006. All spacecraft systems will still see a total of 30 V 
across the rails and no spacecraft performance issues are expected as 
a result. No corrective action is required.

Monday, May 2 (DOY 123)

The monthly teleconference with the Project scientists was held 
today.  Topics of discussion covered instrument team reports on 
current activities including the MAPS workshop, recent Ion and 
Neutral Mass Spectrometer (INMS), Radio and Plasma Wave Science 
(RPWS) and Magnetospheric Imaging Instrument (MIMI) team meetings, 
and media activities surrounding the publication of recent and 
upcoming journal articles of interest.

Tuesday, May 3 (DOY 124)

In this week's science observations, the Magnetometer (MAG) performed 
an 8 hour calibration with the spacecraft rolling about its X-axis. 
The Cosmic Dust Analyzer (CDA) performed a couple of 15 hour 
interstellar dust observations, followed by one 37.5 hour 
interstellar dust observation.  The Visual and Infrared Mapping 
Spectrometer (VIMS) and the Composite Infrared Spectrometer (CIRS) 
made a long 15 hour observation of the E and G rings to get grain 
size measurements. The Imaging Science Subsystem (ISS) performed 
another observation in its Satellite Orbit Campaign, making 
measurements of Methone, Telesto, Janus, Calypso, Prometheus, Pandora 
and Polydeuces.  Ending this week's observations was the Cassini 
Plasma Spectrometer (CAPS) team performing a 13 hour low latitude 
observation to map out compositional and temporal variations in the 
magnetosphere.

In the last week, twelve VIMS cubes were generated and distributed. 
Since Approach Science began in Jan. 2004, 118,330 VIMS cubes and 
243,371 ISS images have been received.  Last week ISS was commanded 
to read out the heap & stack memory; the data will be used to help 
explain the problem causing the instrument to warm-start during 
Instrument Expanded Block (IEB) loads.

Visit the JPL Cassini home page for more information about the 
Cassini Project: http://saturn.jpl.nasa.gov/>

--- Synchronet 3.14a-Win32 NewsLink 1.85
--- SBBSecho 2.11-Win32
* Origin: Internetking BBS Telnet to bbs.hulds.com (1:138/392)
SEEN-BY: 10/1 11/200 331 14/400 34/999 123/500 128/2 187 140/1 222/2 226/0
SEEN-BY: 230/150 249/303 250/306 261/20 38 100 1381 1404 1418 266/1413 275/91
SEEN-BY: 280/1027 320/119 393/68 396/45 633/104 260 267 280 712/848 800/432
SEEN-BY: 801/161 189 2320/105 5030/1256
@PATH: 138/392 250/100 123/500 261/38 633/260 267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.